# SpecEval: Evaluating Code Comprehension in Large Language Models via Program Specifications

### v1

@@ 3. Methodology
In this section, we aim to present a detailed introduction to our framework, SpecEval. We first provide an overview of SpecEval, then describe the evaluation tasks in our framework. For each task, we further introduce the counterfactual analysis, framing code comprehension as the challenge of assessing how controlled input code modifications, through semantic-preserving perturbations, lead to changes in the model’s output. Moreover, for each type of perturbation, we perform Progressive Consistency Analysis to evaluate the consistency of LLM performance across tasks of increasing difficulty. This framework enables a comprehensive assessment of the code comprehension capabilities of LLMs, ranging from superficial to deeper levels of understanding.
3.1. Overview
Fig. 2 demonstrates the overview of our framework, which is characterized by employing formalized program specifications to represent program semantics. We generally adopt the typical architecture of existing black-box evaluations (Chen et al., 2024; Gu et al., 2024), where several evaluation tasks are assigned to the LLMs. The ability of LLMs is evaluated by assessing their performance in these tasks. In our frameworks, four tasks related to program specifications are designed, including Specification Correctness Judgement, Specification Candidates Selection, Specification Infilling, and Specification Generation. For each task, we design five semantic-preserving perturbations performed on the programs and the corresponding ground-truth specifications to create variants with equivalent semantics. Counterfactual analysis is conducted based on the original and mutants, allowing us to study the performance variance of LLMs against semantic-preserving perturbations. Progressive consistency analysis is performed for each type of perturbation to evaluate the consistency of LLM performance on a series of tasks with sequential dependencies in between.
Figure 2. Overview of the evaluation framework SpecEval.
@@ 3.2. Evaluation Tasks and Metrics
We designed four tasks related to formal program specifications. These tasks not only request LLMs to summarize abstract semantic information into formalized statements but also require them to comprehend the given specifications and take appropriate actions based on their understanding.
@@ 3.2.1. Specification Correctness Judgement
Specifications possess correctness, defined by their consistency with program behaviors. A correct specification is about constraints that always hold for all program states in its scope. We denote the actual correctness of the candidate specification $s$ as a function $\mathbf{Cr}:S\rightarrow\{\mathbf{true},\mathbf{false}\}$, which takes a value of $\mathbf{Cr}(s)=\mathbf{true}$ if $s$ is correct and $\mathbf{Cr}(s)=\mathbf{false}$ otherwise. In this task, subject LLM is exploited to decide the correctness of a given candidate specification on a target program. The correct candidates are extracted from the ground-truth specifications of the target program, and the incorrect candidates are obtained by modifying essential components within the ground-truth specification, including replacing variable names with other variables in the same scope, altering variable operators, and swapping statement predicates (i.e., between $\forall$ and $\exists$). Each component of interest is decided randomly whether it will be modified, with equal probabilities (at 50%) of being modified or staying unchanged. At least one modification is performed to ensure the specification is not exactly the same.
@@
    where $\mathbb{I}:\{\mathbf{true},\mathbf{false}\}\rightarrow\{0,1\}$ is the indicator function that returns 1 if and only if the given boolean expression takes a value of $true$.
@@ 3.2.2. Specification Candidates Selection
Specifications possess varying degrees of strength. While weaker specifications such as ensures true can still pass the verification and should be considered correct, they often describe only trivial information about the program. In contrast, stronger specifications provide a more comprehensive expression of the program’s semantics. Therefore, in this task, we consider both the correctness and strength when constructing candidates for the given target program. In particular, four candidate specifications are provided to subject LLM, which is required to choose the most appropriate (i.e., correct and strongest) one for the given target program. The most appropriate candidate (denoted as $\hat{s}$) in four candidates is extracted from the ground-truth specifications for the target program. Other candidates have an equal chance of being modified specifications or trivial specifications. The former is obtained by modifying the ground-truth specification with the same logic mentioned in Section 3.2.1. The latter is randomly chosen from a set of pre-defined simple specifications that are too weak to be incorrect, such as ensures true.
  • Task Formalism. Given a target program $P$, and a set of four candidate specifications $S=\{s_{1},s_{2},s_{3},s_{4}\}$, a selection task $t\in T$ is a 2-tuple $(P,S)$, where the subject model $\mathcal{M}:T\rightarrow S$ is required to return the most appropriate candidate within $S$ for the target program.
@@ 3.2.3. Specification Infilling
In this task, an uncompleted specification (partially masked with a placeholder) is provided to the subject LLM, which is required to fill in the placeholder with appropriate sub-expressions according to the given target program so that the final specification can pass the verifier. The uncompleted specification is acquired by randomly masking specific components of the ground-truth specifications with placeholders. At most, one placeholder is inserted once. The components of interest include array index, variable and method names, and boundary constraints of $\forall$ and $\exists$ statements, which can be identified by parsing specifications into Abstract Syntax Trees.
  • Task Formalism. Given a target program $P$, and an uncompleted specification $s$, an infilling task $t\in T$ is a 2-tuple $(P,s)$, where the subject model $\mathcal{M}:T\rightarrow S$ is required to fill in the placeholder of $s$ according to the target program and return the filled-in version of $s$.
  • Evaluation Metrics. In this task, we use #Pass and Accuracy for measurement. #Pass denotes the number of programs $P$ for which the specification infilled by model $\mathcal{M}$ are correct. Accuracy measures the percentage of correctly infilled specifications. Denoting the correctness of specification $s$ as $\mathbf{Cr}(s)$, the Accuracy for this task can be calculated as:
@@ 3.2.4. Specification Generation
In this task, the subject LLM is provided with a program $P$ with no instrumented specifications. The LLM is exploited to generate a set of specifications $S$ from scratch, describing the general behaviors and functionalities of the target program as closely as possible. The generated specifications are required to include pre/post-conditions for all methods and loop invariants for all loops involved in the target program.
  • Task Formalism. Given a target program $P\in\mathbb{P}$, the subject model $\mathcal{M}:\mathbb{P}\rightarrow 2^{S}$ is required to generate an appropriate set of specifications $S$ for $P$.
  • Evaluation Metrics. In this task, Precision and Recall are adopted to measure the performance of LLMs. Precision measures the percentage of correct specifications that can pass the verifier over all the generated specifications for the target program $P$. Recall measures the percentage of correctly generated ground-truth specifications (i.e., correct and strongest) over all the ground-truth specifications for the target program $P$.
    $Precision(P)=\frac{\sum_{s\in\mathcal{M}(P)}\mathbb{I}(\mathbf{Cr}(s))}{|\mathcal{M}(P)|}\quad Recall(P)=\frac{\sum_{s\in\mathcal{M}(P)}\mathbb{I}(s\in\hat{S}_{P})}{|\hat{S}_{P}|}$
    where $\hat{S}_{P}$ denotes the set of all ground-truth specifications for program $P$. Correspondingly, the average precision and recall on the set $\mathbb{P}$ of all programs in the benchmark can be calculated as:
    $AvgPrecision=\frac{1}{|\mathbb{P}|}\sum\limits_{P\in\mathbb{P}}Precision(P)\quad AvgRecall=\frac{1}{|\mathbb{P}|}\sum\limits_{P\in\mathbb{P}}Recall(P)$
    Apart from precision and recall, we also use #Pass metric to denote the number of programs $P$ for which all specifications generated by model $\mathcal{M}$ can pass the verifier.
@@ 3.3. Counterfactual Analysis
Counterfactual Analysis is a specific technique conducted by observing the variance in the model performance after changing the model inputs in a particular way (Verma et al., 2020). In SpecEval, for each kind of the designed evaluation task in Section 3.2, we respectively adopt some typical semantic-preserving perturbations to analyze the performance of LLMs on the task.
@@ 3.3.1. Semantic-preserving Perturbations
To eliminate potential bias, the perturbations should be performed while adhering to the following criteria.
  • Consistency with the original program. The perturbations should not alter the overall behaviors and functionalities of the program. The perturbed program and the original program should perform exactly the same behaviors if treated as black boxes.
  • Consistency with the original specifications. All specifications for the original program should also be adaptable to the perturbed program without significant structural or semantic modifications. Trivial modifications that have no semantic effects are allowed, such as altering the variable names involved according to specific patterns. The perturbation should not reduce the program points of interest (e.g. methods and loops) where specifications should be generated.
  • Consistency with the evaluation tasks. The perturbed programs should still be adaptable to all the tasks designed for the original program. The perturbations should not alter the problems or the corresponding answers (i.e., ground truth) involved in the tasks. Similar to the previous rule, trivial modifications such as the variable name modification can be performed on the prompts engaged in the tasks if necessary.
@@
Following prior research (Hooda et al., 2024), we adopted five different types of semantic-preserving perturbations that meet the criteria above, including Def-use Break, If-else Flip, Independent Swap, Name random, and Name shuffle.
  • Def-use Break. Def-use chain refers to the relationship between the definitions of a certain variable (where it is assigned with a value, e.g., x = 0) and its subsequent uses (where its value is accessed, e.g., func(x)). The Def-use Break perturbation aims to break the def-use chains within target programs. This is done by assigning the value of a variable var1 to a newly introduced variable var2 and altering all subsequent uses of var1 to var2. In our implementations, the variable name for var2 consists of five random letters or digits. For instance, in the program shown in Fig. 3, a new variable PZYcz inherits the value of variable num1, and replaces the latter in the remaining part of the program. Similar modifications are performed on variable num2 and local variable temp. Since the newly introduced variable completely supersedes the role of the original variable, the semantic-preserving feature of this perturbation is guaranteed.
  • If-else Flip. This perturbation swaps the branches of an if-else statement. This is done by replacing the content of the two branches with each other and negating the branch condition. Obviously, such perturbation does not affect the semantics of the original program.
@@
It is worth noting that for some of the perturbations, the specifications involved in the original program should be mutated simultaneously. For instance, Name Random and Name Shuffle alter all the variable names in the original program, where the specifications should be mutated following the same pattern since the variable names are involved in the specifications as well. If-else Flip perturbation may change the positions of some specifications within the target if-else branch, so the specifications involved need to be migrated. Such modifications to specifications pose no impact on their semantics. Additionally, task prompts involving specifications have to be mutated as well, keeping them equivalent to the original tasks, such as the candidate specifications used in the prompt in Specification Correctness Judgement and Specification Candidate Selection tasks.
@@ 3.3.2. Evaluation Metrics
We adopt different metrics for different tasks for counterfactual analysis. For the task of Judgement, Selection, and Infilling, we adopt Jaccard Distance to measure the impact of perturbations on the task as $J=1-|\mathbb{P}\cap\mathbb{P}^{\prime}|/|\mathbb{P}\cup\mathbb{P}^{\prime}|$ , where $\mathbb{P}$ and $\mathbb{P^{\prime}}$ refer respectively to the set of original and perturbed programs that an LLM can successfully handle for one kind of perturbation in Section 3.3.1. The programs in $\mathbb{P}\cap\mathbb{P}^{\prime}$ refer to the LLM successfully handles both the original and the perturbed versions, indicating that its performance is consistent on these programs, without being interfered by the perturbation, whereas other programs in $\mathbb{P}\cup\mathbb{P}^{\prime}$ witness a varied performance on different versions. The Jaccard Distance measures the impact of perturbations by calculating the percentage of inconsistently-performing programs. A high distance indicates a strong impact that the corresponding perturbation poses.
@@
For the task of Generation, since it is quantitatively measured by precision and recall, we define Average Variance to measure the impact of perturbations for this task. For a specific type of perturbation in Section 3.3.1, the Average Variance of metric $m$ where $m\in\{Precision,Recall\}$ can be calculated as follows:
$v_{m}=\frac{1}{|\mathbb{P}|}\sum\limits_{(P,P^{\prime})\in\mathbb{P}}|m(P^{\prime})-m(P)|\ \quad where\quad\ \mathbb{P}=\{(P,P^{\prime})\ |\ m(P)>0\vee m(P^{\prime})>0\}$
@@
$(P,P^{\prime})$ denotes pairs of the original program and the perturbed program. Note that we only consider $(P,P^{\prime})$ where either $P$ or $P^{\prime}$ achieve a non-zero value of $m$. The average variance of precision and recall is denoted as $v_{prec}$ and $v_{rec}$, respectively.
3.4. Progressive Consistency Analysis
Progressive Consistency describes the degree to which a model can sustain its performance across a sequence of interrelated tasks. (Chen et al., 2024) Intuitively, for a series of sequentially related tasks, models that fail with the preliminary tasks should tend to fail the advanced follow-up tasks as well. The essence of progressive consistency lies in the sequential dependencies within the series of tasks. The completion of follow-up tasks should partially rely on the knowledge obtained from previous tasks or the ability to complete previous tasks. Given a series of tasks designed with the criterion above, we can evaluate the performance consistency between these tasks to assess whether the models comprehend the code progressively.
3.4.1. Sequential Dependencies Between Evaluation Tasks
Concerning the four evaluation tasks described in Section 3.2, the sequential dependency in between mainly lies in the ability to complete them. Each follow-up task relies partially on the ability to complete its previous task.
  • From Judgement to Selection. The selection task essentially requires models to judge the correctness and strength of all given candidates respectively and choose the correct and strongest one based on the judgments.
  • From Selection to Infilling. To complete an infilling task, the models have to comprehend the specification to fill in, analyze the semantics of the missing parts, select the most appropriate code patterns in the program, and convert them to formalized specifications.
  • From Infilling to Generation. Essentially, the specification generation tasks require multiple rounds of infilling without additional contexts for completion. Generating strong specifications requires a thorough understanding of LLMs concerning the semantics of the target program.
With the intuitions above, we can confirm the sequential dependency existing between the evaluation tasks in SpecEval.
3.4.2. Evaluation Metrics
To quantitatively evaluate the consistency described above, we design a specialized metric, Progressive Consistency Score, or PCS for short, for evaluation. The core idea of PCS is to measure how many tasks in a row the model can handle. Due to the sequential dependency described in Section 3.4.1, we expect the models to handle tasks consecutively given that their semantics learning abilities towards sequential tasks with progressive difficulties are leveraged in a progressive manner as well. We denote the results of the four tasks concerning the same program $P$ as a sequence $R(P)=[r_{1},r_{2},r_{3},r_{4}]$, where $r_{i}\in\{\mathbf{true},\mathbf{false}\}$ denotes whether the model succeeds in the $i$-th task. For Judgement, Selection, and Infilling, the tasks are considered successful if the answer equals the ground truth or passes verification. For Generation, we adopt softened criteria and view those trials with $Precision(P)>0$ as successful generations since positive precision indicates some correct specifications (that can pass the verifier) generated and partial semantics comprehended by LLM. Specifically, the PCS for program $P$ (denoted as $\textbf{PCS}(P)$) can be calculated as follows:
  • If the model successfully handles all four tasks, i.e., $R(P)=[\mathbf{true},\mathbf{true},\mathbf{true},\mathbf{true}]$, it receives a full score of $\textbf{PCS}(P)=1$.
  • If the model handles three tasks in a row, i.e., $R(P)=[\mathbf{true},\mathbf{true},\mathbf{true},\mathbf{false}]$ or
    $R(P)=[\mathbf{false},\mathbf{true},\mathbf{true},\mathbf{true}]$, it receives a half score of $\textbf{PCS}(P)=0.5$.
  • If the model handles two tasks in a row, for instance, $R(P)=[\mathbf{true},\mathbf{true},\mathbf{false},\mathbf{false}]$ or
    $R(P)=[\mathbf{true},\mathbf{false},\mathbf{true},\mathbf{true}]$, it receives a quarter score of $\textbf{PCS}(P)=0.25$.
  • If the model can only handle no consecutive tasks such as $R(P)=[\mathbf{true},\mathbf{false},\mathbf{false},\mathbf{true}]$ or fails to handle any task, we recognize that no consistency is shown between the tasks, and it receives a score of $\textbf{PCS}(P)=0$.
Correspondingly, the average Progressive Consistency Score for the set $\mathbb{P}$ of all programs in the benchmark can be calculated as $\mathbf{PCS}=\frac{1}{|\mathbb{P}|}\sum_{P\in\mathbb{P}}\mathbf{PCS}(P)$. It can be observed that PCS is defined based on the number of consecutive tasks handled by the model. Higher scores indicate more tasks consecutively handled and intuitively higher progressive consistency in the model’s performance. Additionally, considering the difficulty of completing all four tasks, the score obtained is designed to grow exponentially from 0.25 to 1 as the number of consecutively handled tasks increases. Compared to the Incremental Consistency Score (ICS) designed by Chen et al. (Chen et al., 2024) to evaluate the reasoning ability concerning program runtime behaviors, PCS considers more potential consistency between all the tasks, especially the follow-up ones. For instance, $R(P)=[\mathbf{false},\mathbf{true},\mathbf{true},\mathbf{true}]$ receives an ICS of 0, whereas the score is 0.5 for PCS, since such performance still demonstrates consistency between the follow-up tasks. We assume that the consistency between the latter tasks is meaningful despite the failure in preliminary tasks and relevant inconsistency.

### v2

@@ 3. Methodology
In this Section, we aim to present a detailed introduction to our framework. We first describe the evaluation tasks in our framework. For each task, we introduce the counterfactual analysis, framing code comprehension as the challenge of assessing how controlled input code modifications, through semantic-preserving perturbations, lead to changes in the model’s output. This framework enables a comprehensive assessment of the code comprehension capabilities of LLMs, ranging from superficial to deeper levels of understanding.
@@ 3.1. Evaluation Tasks and Metrics
We designed four tasks related to formal program specifications. These tasks not only request LLMs to summarize abstract semantic information into formalized statements but also require them to comprehend the given specifications and take appropriate actions based on their understanding. Some tasks directly correspond to specific software engineering techniques, while others implicitly assess the LLMs’ ability to comprehend formal languages.
@@ 3.1.1. Specification Correctness Judgement
Specifications possess correctness, defined by their consistency with program behaviors. A correct specification is about constraints that always hold for all program states in its scope. We denote the actual correctness of the candidate specification $s$ as a function $\mathbf{Cr}:S\rightarrow\{\mathbf{true},\mathbf{false}\}$, which takes a value of $\mathbf{Cr}(s)=\mathbf{true}$ if $s$ is correct and $\mathbf{Cr}(s)=\mathbf{false}$ otherwise. In this task, subject LLM is exploited to decide the correctness of a given candidate specification on a target program. This task closely aligns with software model checking (Jhala and Majumdar, 2009), a static technique to prove whether a program satisfies specific properties. The correct candidates are extracted from the ground-truth specifications of the target program. The incorrect candidates are obtained by modifying essential components within the ground-truth specification, including replacing variable names with other variables in the same scope, altering variable operators, and swapping statement predicates (i.e., between $\forall$ and $\exists$). Each component of interest is decided randomly whether it will be modified, with equal probabilities (at 50%) of being modified or staying unchanged. At least one modification is performed to ensure the specification does not stay the same. Runtime checking is performed on the mutated candidates to ensure their incorrectness. A candidate should be mutated again if it remains correct after mutation.
@@
    where $\mathbb{I}:\{\mathbf{true},\mathbf{false}\}\rightarrow\{0,1\}$ is the indicator function that maps $\mathbf{true}$ to 1 and $\mathbf{false}$ to 0.
@@ 3.1.2. Specification Candidates Selection
[…]
  • Task Formalism. Given a target program $P$ and a set of four candidate specifications $S=\{s_{1},s_{2},s_{3},s_{4}\}$, a selection task $t\in T$ is a 2-tuple $(P,S)$, where the subject model $\mathcal{M}:T\rightarrow S$ should return the most appropriate candidate within $S$ for $P$. The ground truth for this task is the correct candidate, denoted as $\hat{s}$.
@@ 3.1.3. Specification Infilling
In this task, an uncompleted specification (partially masked with a placeholder) is provided to the subject LLM, which is required to fill in the placeholder with appropriate sub-expressions according to the given target program so that the final specification can pass the verifier. The uncompleted specification is acquired by randomly masking specific components of the ground-truth specifications with placeholders. Only one placeholder is inserted for each task. The components of interest include array index, variable and method names, and boundary constraints of $\forall$ and $\exists$ statements, which can be identified by parsing specifications into Abstract Syntax Trees. The task formulation is highly similar to code infilling (Fried et al., 2022), a derived coding task widely adopted in various LLM benchmarks (Bavarian et al., 2022; Gong et al., 2024), except that it is the specification that is to be infilled here rather than the code.
  • Task Formalism. Given a target program $P$, and an uncompleted specification $s$, an infilling task $t\in T$ is a 2-tuple $(P,s)$, where the subject model $\mathcal{M}:T\rightarrow S$ is required to fill in the placeholder of $s$ according to the target program and return the filled-in version.
  • Evaluation Metrics. In this task, we use #Pass and Accuracy for measurement. #Pass denotes the number of programs $P$ for which the specification infilled by model $\mathcal{M}$ are correct. Accuracy measures the percentage of correctly infilled specifications. The Accuracy for this task can be calculated as:
@@ 3.1.4. Specification Generation
In this task, the subject LLM is provided with a program $P$ with no instrumented specifications. The LLM is exploited to generate a set of specifications $S$ from scratch, describing the general behaviors and functionalities of the target program as closely as possible. The generated specifications should include pre/post-conditions for all methods and loop invariants for all loops involved in the target program. Such task formulation is frequently adopted in recent works on program specification generation techniques (Ma et al., 2024; Wen et al., 2024).
  • Task Formalism. Given a program $P$, the model $\mathcal{M}:\mathbb{P}\rightarrow 2^{S}$ should generate an appropriate set of specifications $S$ for $P$.
  • Evaluation Metrics. In this task, Precision and Recall are adopted to measure the performance of LLMs. Precision measures the percentage of correct specifications that can pass the verifier over all the generated specifications. Recall measures the percentage of correctly generated ground-truth specifications (i.e., correct and strongest) over all the ground-truth specifications for the program $P$. The above metrics can be calculated by
    $Prec(P)=\frac{\sum_{s\in\mathcal{M}(P)}\mathbb{I}(\mathbf{Cr}(s))}{|\mathcal{M}(P)|}\quad Rec(P)=\frac{\sum_{\hat{s}\in\hat{S}_{P}}\mathbb{I}(\hat{s}\in\mathcal{M}(P))}{|\hat{S}_{P}|}$
    where $\hat{S}_{P}$ denotes the set of all ground-truth specifications for program $P$. Apart from precision and recall, we also use #Pass metric to denote the number of programs $P$ for which all specifications generated by model $\mathcal{M}$ can pass the verifier.
It is noteworthy that calculating Recall necessitates determining whether each ground truth specification is covered by the model’s outputs, i.e., whether $\hat{s}\in\mathcal{M}(P)$ holds. This involves determining the equivalence between a ground truth specification $\hat{s}\in\hat{S}_{P}$ and each of the model-generated specifications $s\in\mathcal{M}(P)$. To address this issue, we utilize a bidirectional implication operator to combine the two specifications into a new composite specification $\hat{s}\Leftrightarrow s$. The equivalence between the two specifications is then determined by evaluating the correctness of this composite specification. Thus, if there exists a model-generated specification $s\in\mathcal{M}(P)$ that is equivalent to $\hat{s}$, we can conclude that $\hat{s}\in\mathcal{M}(P)$ and vice versa.
@@ 3.2. Counterfactual Analysis
Counterfactual Analysis is a technique conducted by observing the variance in the model performance after changing the model inputs in a particular way (Verma et al., 2020). In SpecEval, we respectively adopt several typical semantic-preserving perturbations to analyze the performance variance of LLMs on each kind of the evaluation tasks.
@@ 3.2.1. Semantic-preserving Perturbations
Among the common semantically equivalent program transformation strategies (Hooda et al., 2024; Le-Cong et al., 2024), we adopted the following five general ones as the perturbations.
  • Def-use Break. Def-use chain refers to the relationship between the definitions of a certain variable (where it is assigned with a value, e.g., x = 0) and its subsequent uses (where its value is accessed, e.g., func(x)). The Def-use Break perturbation aims to break these chains within target programs by assigning the value of a variable var1 to a newly added variable var2 and altering all subsequent uses of var1 to var2. In our implementations, the variable name for var2 consists of five random letters or digits. For instance, in the program shown in Fig. 2, a new variable PZYcz inherits the value of variable num1, and replaces the latter in the remaining part of the program. Similar modifications are performed on variable num2 and local variable temp. Since the new variable completely supersedes the role of the original variable, the perturbation is guaranteed to be semantic-preserving.
  • If-else Flip. This perturbation swaps the branches of an if-else statement. This is done by replacing the content of the two branches with each other and negating the branch condition. Such perturbation does not affect the program semantics.
@@
It is worth noting that for some of the perturbations, the specifications involved in the original program should be mutated simultaneously. For instance, Name Random and Name Shuffle alter all the variable names in the original program, where the specifications should undergo identical lexical modifications since the variable names are also involved in the specifications. If-else Flip perturbation may change the positions of some specifications within the target if-else branch, so the specifications involved need to be migrated. Such modifications to specifications pose no impact on their semantics. Furthermore, task prompts involving specifications must also be mutated to keep them equivalent to the original tasks, such as the candidate specifications in Judgement and Selection tasks.
@@ 3.2.2. Evaluation Metrics
We adopt different metrics for different tasks in counterfactual analysis. For Judgement, Selection, and Infilling, we adopt Jaccard Distance to measure the impact of perturbations on the task as $J=1-|\mathbb{P}\cap\mathbb{P}^{\prime}|/|\mathbb{P}\cup\mathbb{P}^{\prime}|$ , where $\mathbb{P}$ and $\mathbb{P^{\prime}}$ refer respectively to the set of original and perturbed programs that an LLM can successfully handle for one kind of perturbation in Section 3.2.1. The programs in $\mathbb{P}\cap\mathbb{P}^{\prime}$ refer to the LLM successfully handles both the original and the perturbed versions, indicating that its performance is consistent on these programs, without being interfered by the perturbation, whereas other programs in $\mathbb{P}\cup\mathbb{P}^{\prime}$ witness a varied performance on different versions. The Jaccard Distance measures the impact of perturbations by calculating the percentage of inconsistently-performing programs. A high distance indicates a strong impact that the corresponding perturbation poses.
@@
For the Generation task, since it is quantitatively measured by precision and recall, we define Average Variance to measure the impact of perturbations for this task. For a specific type of perturbation in Section 3.2.1, the Average Variance of metric $m$ where $m\in\{Precision,Recall\}$ can be calculated as
$v_{m}=\frac{1}{|\mathbb{P}|}\sum\limits_{(P,P^{\prime})\in\mathbb{P}}|m(P^{\prime})-m(P)|$
@@
where =$\mathbb{P}=\{(P,P^{\prime})\ |\ m(P)>0\vee m(P^{\prime})>0\}$, $(P,P^{\prime})$ denotes pairs of the original and the perturbed programs. To keep consistent with the Jaccard Distance, we only consider $(P,P^{\prime})$ where either $P$ or $P^{\prime}$ achieve a non-zero value of $m$. The average variance of precision and recall is denoted as $v_{prec}$ and $v_{rec}$, respectively.